5. Let X represents the number of computers in an Australian household, for those that own a
computer.

a. Find and interpret the expected number of computers in a randomly selected Australian
household. (2 marks)
b. What is the probability that a randomly selected Australian household will have more than 2
computers?
c. Find V (4X +2). (1 mark)
d. Find E(30X + 20]. (1 mark)
1
Expert's answer
2020-04-30T18:36:15-0400

"\begin{matrix}
 x & 1 & 2 & 3 & 4 & 5 \\
 p(x) & 0.25 & 0.33 & 0.17 & 0.15 & 0.10
\end{matrix}"


a. The expected number of computers in a randomly selected Australian household


"E(X)=1(0.25)+2(0.33)+3(0.17)+4(0.15)+5(0.10)=""=2.52"

The expexted number of computers in an Australian household, for those that own a computer is

"2.52" computers.


b. The probability that a randomly selected Australian household will have more than 2 computers is


"P(X>2)=0.17+0.15+0.10=0.42"

c.


"V(X)=(1-2.52)^2(0.25)+(2-2.52)^2(0.33)+""+(3-2.52)^2(0.17)+(4-2.52)^2(0.15)+""+(5-2.52)^2(0.10)=1.6496"

"V(4X+2)=4^2\cdot V(X)=16\cdot1.6496=26.3936"

d.


"E(30X+20)=30E(X)+20=30\cdot2.52+20=95.6"